Nnesha (nnesha):

help please.......................................... Find the local maximum of the function y = x3 + 3x2 - 2x + 4

Nnesha (nnesha):

ct one: a. (-2.3, 12.3) b. (-3.8, 0) c. (0, 4) d. (0.3, 3.7)

OpenStudy (amoodarya):

find y' then solve y'=0 then determine the sign of y' then find local max

Nnesha (nnesha):

i didn't understand :(

OpenStudy (amoodarya):

\[y=x^3+3x^2-2x+4\\y'=3x^2+6x-2\\y'=0\\3x^2+6x-2=0\\\]

Nnesha (nnesha):

i got a is that right

OpenStudy (anonymous):

a. is correct. The exact symbolic answer is:\[\left\{\frac{1}{3} \left(-3-\sqrt{15}\right),\frac{10 \sqrt{\frac{5}{3}}}{3}+8\right\} \]